# Tony Lindeberg

### v5

@@ 3.1 Generalized idealized models of complex cells
The motivation for adding derivatives of higher order than 2 are: (i) to enable more narrow orientation selectivity properties and (ii) in Lindeberg (2024a) it was shown that models of simple cells up to order 4 lead to better agreement with the orientation selectivity histogram of simple cells recorded by Goris et al. (2015) than simple cells up to order 2.
@@
$L_{\varphi\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=\\ =T_{\varphi,\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})*f(x_{1},x_{2}),$ (9)
$L_{\varphi\varphi\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=\\ =T_{\varphi\varphi,\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})*f(x_{1},x_{2}),$ (10)
@@
${\cal Q}_{\varphi,12,\text{int}}L=\\ =\sqrt{\sum_{m\in\{1,2\}}g(\cdot,\cdot;\;\gamma^{2}\Sigma_{\varphi})*C_{\varphi}^{m-1}\,L_{\varphi^{m},\scriptsize\mbox{norm}}^{2}(\cdot,\cdot)}$ (11)
${\cal Q}_{\varphi,1234,\text{int}}L=\\ =\sqrt{\sum_{m\in\{1,2,3,4\}}g(\cdot,\cdot;\;\gamma^{2}\Sigma_{\varphi})*C_{\varphi}^{m-1}\,L_{\varphi^{m},\scriptsize\mbox{norm}}^{2}(\cdot,\cdot)}$ (12)
${\cal Q}_{\varphi,34,\text{int}}L=\\ =\sqrt{\sum_{m\in\{3,4\}}g(\cdot,\cdot;\;\gamma^{2}\Sigma_{\varphi})*C_{\varphi}^{m-3}\,L_{\varphi^{m},\scriptsize\mbox{norm}}^{2}(\cdot,\cdot)},$ (13)
@@
${\cal Q}_{\varphi,12,\text{int}}L$ ${\cal Q}_{\varphi,1234,\text{int}}L$ ${\cal Q}_{\varphi,34,\text{int}}L$
$\kappa=1$ Refer to caption Refer to caption Refer to caption
$\kappa=2$ Refer to caption Refer to caption Refer to caption
$\kappa=4$ Refer to caption Refer to caption Refer to caption
$\kappa=8$ Refer to caption Refer to caption Refer to caption
Figure 3: Graphs of the orientation selectivity for the generalized integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$ according to (11), ${\cal Q}_{\varphi,1234,\text{int}}L$ according to (12) and ${\cal Q}_{\varphi,34,\text{int}}L$ according to (13), which combine integrated squared values of affine Gaussian derivative responses for different combinations of orders of integration $m\in\{1,2,3,4\}$ when applied to an ideal sine wave of the form (6) with the angular frequency of the sine wave adapted so as to evoke a maximally strong response over the angular frequencies according to (22), (23) and (24), respectively. The resulting orientation selectivity curves are shown for different values of the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$, which parameterizes the degree of elongation of the receptive fields. (top row) Results for $\kappa=1$. (second row) Results for $\kappa=2$. (third row) Results for $\kappa=4$. (bottom row) Results for $\kappa=8$. (Horizontal axes: orientation $\theta\in[-\pi/2,\pi/2]$. Vertical axes: Amplitude of the receptive field response relative to the maximum response obtained for $\theta=0$.)
3.2 Orientation selectivity curves for the generalized idealized models of complex cells
To characterize the orientation selectivity properties for the generalized integrated affine quasi quadrature measures (11), (12) and (13), let us first compute the responses $L_{\varphi^{m},\scriptsize\mbox{norm}}$ for the underlying models of simple cells $T_{\varphi^{m},\scriptsize\mbox{norm}}$ to a sine wave (6) according to Equations (29) and (35) in Lindeberg (2025a)
@@ 3.2 Orientation selectivity curves for the generalized idealized models of complex cells
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}T_{0,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2};\;\sigma_{1},\sigma_{2})\end{split}$
[…]
$\begin{split}&=\omega\,\sigma_{1}\cos(\theta)\,e^{-\frac{1}{2}\omega^{2}(\sigma_{1}^{2}\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta)}\end{split}$
×$\begin{split}&\phantom{==}\times\cos(\omega\cos(\theta)\,x_{1}+\omega\sin(\theta)\,x_{2}+\beta),\end{split}$ (14)
@@
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}T_{00,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2};\;\sigma_{1},\sigma_{2})\end{split}$
[…]
$\begin{split}&=-\omega^{2}\,\sigma_{1}^{2}\cos^{2}(\theta)\,e^{-\frac{1}{2}\omega^{2}(\sigma_{1}^{2}\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta)}\end{split}$
×$\begin{split}&\phantom{==}\times\sin(\omega\cos(\theta)\,x_{1}+\omega\sin(\theta)\,x_{2}+\beta),\end{split}$ (15)
and according to Equations (36) and (42) in Lindeberg (2024a)
@@
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}T_{000,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2};\;\sigma_{1},\sigma_{2})\end{split}$
[…]
$\begin{split}&=-\omega^{3}\,\sigma_{1}^{3}\cos^{3}(\theta)\,e^{-\frac{1}{2}\omega^{2}(\sigma_{1}^{2}\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta)}\end{split}$
×$\begin{split}&\phantom{==}\times\cos(\omega\cos(\theta)\,x_{1}+\omega\sin(\theta)\,x_{2}+\beta),\end{split}$ (16)
@@
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}T_{0000,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2};\;\sigma_{1},\sigma_{2})\end{split}$
[…]
$\begin{split}&=\omega^{4}\,\sigma_{1}^{4}\cos^{4}(\theta)\,e^{-\frac{1}{2}\omega^{2}(\sigma_{1}^{2}\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta)}\end{split}$
×$\begin{split}&\phantom{==}\times\sin(\omega\cos(\theta)\,x_{1}+\omega\sin(\theta)\,x_{2}+\beta),\end{split}$ (17)
@@
Let us next integrate the squares of these expressions spatially using a Gaussian window function with relative integration scale $\gamma>1$, with the actual calculations performed in Mathematica, and leading to results that are unfortunately too complex to be reproduced here.
Let us furthermore for the parameterization of the vertical scale parameter $\sigma_{2}=\kappa\,\sigma_{1}$ consider the angular frequencies $\hat{\omega_{1}}$, $\hat{\omega_{2}}$, $\hat{\omega_{3}}$ and $\hat{\omega_{4}}$ for which these expressions assume their maxima over angular frequencies according to Equations (33) and (38) in Lindeberg (2025a)
@@
and according to Equations (40) and (45) in Lindeberg (2024a)
@@
Figure 3 shows the resulting orientation selectivity curves that we then obtain for different values of the scale parameter ratio $\kappa$, when using the relative integration scale $\gamma=1/\sqrt{2}$ for the spatial integration stage and the weighting factor $C=1/\sqrt{2}$ when combining the responses for different orders $m$ of differentiation. As can be seen from these graphs, for all the three generalized integrated affine quasi quadrature measures, the orientation selectivity curves become sharper with increasing values of the scale parameter ratio $\kappa$. In this respect, these results are consistent with the previously reported results in Lindeberg (2025a; 2024a).
@@
Specifically, in relation to the previously recorded orientation selectivity curves for biological neurons reported in Nauhaus et al. (2008), and which span a variability in orientation selectivity from wide to narrow orientation selectivity properties, these results are consistent with what would be the case if the receptive fields in the primary visual cortex would span a variability in the degree of elongation, as proposed as a working hypothesis in Lindeberg (2023) Section 3.2.1 and further investigated in Lindeberg (2024a).
Biological complex cells (n = 184) Histogram over $|R|$ for ${\cal Q}_{\varphi,12,\text{pt}}L$
Refer to caption Refer to caption
Graph of resultant $|R(\kappa)|$ for ${\cal Q}_{\varphi,12,\text{int}}L$ Graph of resultant $|R(\kappa)|$ for ${\cal Q}_{\varphi,1234,\text{int}}L$ Graph of resultant $|R(\kappa)|$ for ${\cal Q}_{\varphi,34,\text{int}}L$
Refer to caption Refer to caption Refer to caption
Histogram over $|R|$ for ${\cal Q}_{\varphi,12,\text{int}}L$ Histogram over $|R|$ for ${\cal Q}_{\varphi,1234,\text{int}}L$ Histogram over $|R|$ for ${\cal Q}_{\varphi,34,\text{int}}L$
Refer to caption Refer to caption Refer to caption
Combined histogram over $|R|$ for ${\cal Q}_{\varphi,12,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$
Figure 4: Orientation selectivity analysis for the resultant of the orientation selectivity curves, with comparison to biological orientation selectivity histograms by Goris et al. (2015). (first row, left) Resultant histogram obtained from neurophysiological recordings of 184 complex cells in Macaque monkeys. (first row, right) Resultant histogram of the pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{pt}}L$ according to (3) based on the previous analysis in (Lindeberg 2024a). (second row) Graphs of the resultant $|R|$ as function of the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ for the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$, ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ according to (11), (12) and (13). (third row) Histograms over the resultant $|R|$ for the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$, ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ according to (11), (12) and (13), when assuming a uniform distribution over a logarithmic parameterization of the scale parameter ratio $\kappa$. (fourth row) Combined histogram of ${\cal Q}_{\varphi,12,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$, when assuming equal numbers of complex cells for these two types. (Horizontal axes in rows 1, 3 and 4: 10 quantized bins over the resultant $|R|\in[0,1]$. Vertical axes in rows 1, 3 and 4: bin counts.) (Horizontal axes in row 2: scale parameter ratio $\kappa$. Vertical axes in row 2: resultant $|R|$.)
@@
3.3 Orientation selectivity histograms for the generalized idealized models of complex cells
@@
To aim at a more quantitative analys, let us compare the result of our idealized integrated models of complex cells with the quantitative measurements of orientation selectivity histograms reported by Goris et al. (2015). They accumulated histograms of the absolute value of the resultant $|R|$ with the underlying complex-valued resultant of an orientation selectivity curve $r(\theta)$ of the form
[…]
@@
The left figure in the top row in Figure 4 gives a schematic illustration of the results that they obtained, reflecting a significant variability in wide vs. narrow orientation selectivity properties for different biological complex cells. The right figure in the top row in Figure 4 shows a corresponding prediction of a histogram of the resultant of the orientation selectivity curves obtained in Lindeberg (2024a), based on the pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{pt}}L$ in (3), based on assuming that the scale parameter ratio $\kappa$ would be uniformly distributed on a logarithmic scale over the interval $[1/\kappa_{\max},\kappa_{max}]$ for the arbitrary choice of $\kappa_{max}=8$.
@@
As can be seen from the comparison between the biological results and the idealized modelling results, the distribution over the resultant $|R|$ is somewhat biased toward both smaller and larger values of $|R|$, compared to the neurophysiologically accumulated resultant histograms. A natural question to ask is hence if this behaviour would be different if using more developed models of complex cells, that also comprise a spatial integration stage and derivatives of higher order than 2.
@@
The second row in Figure 4 show the result of computing the resultant measure $|R|$ as a function of the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ for each one of the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$, ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ according to (11), (12) and (13).
@@
The third row in Figure 4 show corresponding resultant histograms for each of these integrated affine quasi quadrature measures, when assuming a uniform distribution over a logarithmic parameterization of the scale parameter ratio $\kappa$ over the interval $[1/\kappa_{\max},\kappa_{\max}]$ for the again arbitrary value of $\kappa_{\max}=8$. Such a logarithmic distribution constitutes a natural default prior for a strictly positive variable according to the (Jaynes 1968). The bottom row in Figure 4 shows a combined histogram of the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$, when assuming an equal number of idealized complex cells for these two types.
@@
As can be seen from these histograms, the use of the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$, ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ leads to resultant histograms with different shapes than for pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{pt}}L$ previously studied in Lindeberg (2024a). Specifically, except for the peak at the bin corresponding to $|R|\in[0.6,0.7]$, the resultant histogram of ${\cal Q}_{\varphi,1234,\text{int}}L$ is more uniform and without any bias the bias towards either smaller values of the resultant $|R|$ towards the bins $|R|\in[0.1,0.2]$ and $|R|\in[0.7,0.8]$ as for the resultant histogram of the pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{pt}}L$.
@@
Additionally, when including third- and fourth-order terms, the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ lead to accumulations to the bins $|R|\in[0.8,0.9]$ and $|R|\in[0.9,1.0]$, while those bins are not really reached by the pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{pt}}L$. In these respects, this analysis suggests that the mechanisms of spatial integration and inclusion of higher-order terms than mere first- and second-order terms may be important to more quantitatively model the orientation selectivity properties of complex cells. Furthermore, one may speculate if a suitable reformulation of the non-linearities in the composed integrated affine quasi quadrature measures ${\cal Q}_{\varphi,1234,\text{int}}L$ could move the peak for the bin $|R|\in[0.6,0.7]$ to the bin positions $|R|\in[0.3,0.4]$ and $|R|\in[0.5,0.6]$ that lead to peaks in the resultant histogram for the biological complex cells.

### v7

@@ Results
The motivation for adding derivatives of higher order than 2 are: (i) to enable more narrow orientation selectivity properties and (ii) in Lindeberg [14] it was shown that models of simple cells up to order 4 lead to better agreement with the orientation selectivity histogram of simple cells recorded by Goris et al. [15] than simple cells up to order 2.
@@
$L_{\varphi\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=T_{\varphi\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})*f(x_{1},x_{2}),$ (9)
$L_{\varphi\varphi\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})=T_{\varphi\varphi\varphi\varphi,\scriptsize\mbox{norm}}(x_{1},x_{2};\;\sigma_{\varphi},\Sigma_{\varphi})*f(x_{1},x_{2}),$ (10)
@@
${\cal Q}_{\varphi,12,\text{int}}L=\sqrt{\sum_{m\in\{1,2\}}g(\cdot,\cdot;\;\gamma^{2}\,\Sigma_{\varphi})*C_{\varphi}^{m-1}\,L_{\varphi^{m},\scriptsize\mbox{norm}}^{2}(\cdot,\cdot;\;\sigma_{\varphi},\Sigma_{\varphi})}$ (11)
${\cal Q}_{\varphi,1234,\text{int}}L=\sqrt{\sum_{m\in\{1,2,3,4\}}g(\cdot,\cdot;\;\gamma^{2}\,\Sigma_{\varphi})*C_{\varphi}^{m-1}\,L_{\varphi^{m},\scriptsize\mbox{norm}}^{2}(\cdot,\cdot;\;\sigma_{\varphi},\Sigma_{\varphi})}$ (12)
${\cal Q}_{\varphi,34,\text{int}}L=\sqrt{\sum_{m\in\{3,4\}}g(\cdot,\cdot;\;\gamma^{2}\,\Sigma_{\varphi})*C_{\varphi}^{m-3}\,L_{\varphi^{m},\scriptsize\mbox{norm}}^{2}(\cdot,\cdot;\;\sigma_{\varphi},\Sigma_{\varphi})},$ (13)
@@
Orientation selectivity curves for the generalized idealized models of complex cells
To characterize the orientation selectivity properties for the generalized integrated affine quasi quadrature measures (11), (12) and (13), let us first compute the responses $L_{\varphi^{m},\scriptsize\mbox{norm}}$ for the underlying models of simple cells $T_{\varphi^{m},\scriptsize\mbox{norm}}$ to a sine wave (6) according to Eqs. (29) and (35) in Lindeberg [16]
@@
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}T_{0,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2};\;\sigma_{1},\sigma_{2})\,f(x_{1}-\xi_{1},x_{2}-\xi_{2})\,d\xi_{1}\,d\xi_{2}\end{split}$
$\begin{split}&=\omega\,\sigma_{1}\cos(\theta)\,e^{-\frac{1}{2}\omega^{2}(\sigma_{1}^{2}\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta)}\,\cos(\omega\cos(\theta)\,x_{1}+\omega\sin(\theta)\,x_{2}+\beta),\end{split}$ (14)
@@
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}T_{00,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2};\;\sigma_{1},\sigma_{2})\,f(x_{1}-\xi_{1},x_{2}-\xi_{2})\,d\xi_{1}\,d\xi_{2}\end{split}$
$\begin{split}&=-\omega^{2}\,\sigma_{1}^{2}\cos^{2}(\theta)\,e^{-\frac{1}{2}\omega^{2}(\sigma_{1}^{2}\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta)}\,\sin(\omega\cos(\theta)\,x_{1}+\omega\sin(\theta)\,x_{2}+\beta),\end{split}$ (15)
and according to Eqs. (36) and (42) in the supplementary material of Lindeberg [14]
@@
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}T_{000,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2};\;\sigma_{1},\sigma_{2})\,f(x_{1}-\xi_{1},x_{2}-\xi_{2})\,d\xi_{1}\,d\xi_{2}\end{split}$
$\begin{split}&=-\omega^{3}\,\sigma_{1}^{3}\cos^{3}(\theta)\,e^{-\frac{1}{2}\omega^{2}(\sigma_{1}^{2}\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta)}\,\cos(\omega\cos(\theta)\,x_{1}+\omega\sin(\theta)\,x_{2}+\beta),\end{split}$ (16)
@@
$\begin{split}&=\int_{\xi_{1}=-\infty}^{\infty}\int_{\xi_{2}=-\infty}^{\infty}T_{0000,\scriptsize\mbox{norm}}(\xi_{1},\xi_{2};\;\sigma_{1},\sigma_{2})\,f(x_{1}-\xi_{1},x_{2}-\xi_{2})\,d\xi_{1}\,d\xi_{2}\end{split}$
$\begin{split}&=\omega^{4}\,\sigma_{1}^{4}\cos^{4}(\theta)\,e^{-\frac{1}{2}\omega^{2}(\sigma_{1}^{2}\cos^{2}\theta+\sigma_{2}^{2}\sin^{2}\theta)}\,\sin(\omega\cos(\theta)\,x_{1}+\omega\sin(\theta)\,x_{2}+\beta),\end{split}$ (17)
@@
Let us next integrate the squares of these expressions spatially using a Gaussian window function with relative integration scale $\gamma>1$, with the actual calculations performed in Wolfram Mathematica and leading to results that are unfortunately too complex to be reproduced here.
Let us furthermore for the parameterization of the vertical scale parameter $\sigma_{2}=\kappa\,\sigma_{1}$ consider the angular frequencies $\hat{\omega_{1}}$, $\hat{\omega_{2}}$, $\hat{\omega_{3}}$ and $\hat{\omega_{4}}$ for which these expressions assume their maxima over angular frequencies according to Eqs. (33) and (38) in Lindeberg [16]
@@
and according to Eqs. (40) and (45) in the supplementary material of Lindeberg [14]
@@
Fig. 3 shows the resulting orientation selectivity curves that we then obtain for different values of the scale parameter ratio $\kappa$, when using the relative integration scale $\gamma=1/\sqrt{2}$ for the spatial integration stage and the weighting factor $C=1/\sqrt{2}$ when combining the responses for different orders $m$ of differentiation. As can be seen from these graphs, for all the three generalized integrated affine quasi quadrature measures, the orientation selectivity curves become sharper with increasing values of the scale parameter ratio $\kappa$. In this respect, these results are consistent with the previously reported results in Lindeberg [16, 14].
${\cal Q}_{\varphi,12,\text{int}}L$
${\cal Q}_{\varphi,1234,\text{int}}L$
${\cal Q}_{\varphi,34,\text{int}}L$
Fig 3: Graphs of the orientation selectivity curves for the generalized integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$ according to (11), ${\cal Q}_{\varphi,1234,\text{int}}L$ according to (12) and ${\cal Q}_{\varphi,34,\text{int}}L$ according to (13), and which combine integrated squared values of affine Gaussian derivative responses for different combinations of orders of integration $m\in\{1,2,3,4\}$ when applied to an ideal sine wave of the form (6) with the angular frequency of the sine wave adapted so as to evoke a maximally strong response over the angular frequencies according to (22), (23) and (24), respectively. The resulting orientation selectivity curves are shown for different values of the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}\in\{1,2,4,8\}$, which parameterizes the degree of elongation of the receptive fields. (Horizontal axes: orientation $\theta\in[-\pi/2,\pi/2]$. Vertical axes: Amplitude of the receptive field response relative to the maximum response obtained for $\theta=0$.)
@@
Specifically, in relation to the previously recorded orientation selectivity curves for biological neurons reported in Nauhaus et al. [22], and which span a variability in orientation selectivity from wide to narrow orientation selectivity properties, these results are consistent with what would be the case if the receptive fields in the primary visual cortex would span a variability in the degree of elongation, as proposed as a working hypothesis in Lindeberg [10] Sec. 3.2.1 and further investigated in Lindeberg [14].
@@
Orientation selectivity histograms for the generalized idealized models of complex cells
@@
To aim at a more quantitative analys, let us compare the result of our idealized integrated models of complex cells with the quantitative measurements of orientation selectivity histograms reported by Goris et al. [15]. They accumulated histograms of the absolute value of the resultant $|R|$ with the underlying complex-valued resultant of an orientation selectivity curve $r(\theta)$ of the form
$R=\frac{\int_{\theta=-\pi}^{\pi}r(\theta)\,e^{2i\theta}d\theta}{\int_{\theta=-\pi}^{\pi}r(\theta)\,d\theta}.$ (25)
@@
Fig. 4a gives a schematic illustration of the results that they obtained, reflecting a significant variability in wide vs. narrow orientation selectivity properties for different biological complex cells. Fig. 4b shows a corresponding prediction of a histogram of the resultant of the orientation selectivity curves obtained in Lindeberg [14], based on the pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{pt}}L$ in (3), and assuming that the scale parameter ratio $\kappa$ would be uniformly distributed on a logarithmic scale over the interval $1/\kappa_{\max},\kappa_{max}]$ for the arbitrary choice of $\kappa_{max}=8$.
(a) Biological complex cells (n = 184)
(b) Histogram over $|R|$ for ${\cal Q}_{\varphi,12,\text{pt}}L$
Fig 4: Orientation selectivity histogram of the resultant of biological complex cells accumulated by Goris et al. [15] with comparison to histogram of the resultant from previous orientation selectivity analysis of maximally simplified models of complex cells in Lindeberg [14]. (Horizontal axes: 10 quantized bins over the resultant $|R|\in[0,1]$. Vertical axes: bin counts.)
@@
As can be seen from the comparison between the biological results and the idealized modelling results, the distribution over the resultant $|R|$ is somewhat biased towards both smaller and larger values of $|R|$, specifically regarding the peaks in the histogram at the bins $|R|\in[0.1,0.2]$ and $|R|\in[0.7,0.8]$, compared to the neurophysiologically accumulated resultant histograms. A natural question to ask is hence if this behaviour would be different if using more developed models of complex cells, that also comprise a spatial integration stage and derivatives of higher order than 2.
@@
Fig. 5 shows the result of computing the resultant measure $|R|$ as a function of the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ for each one of the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$, ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ according to (11), (12) and (13).
(a) Graph of resultant $|R(\kappa)|$ for ${\cal Q}_{\varphi,12,\text{int}}L$
(b) Graph of resultant $|R(\kappa)|$ for ${\cal Q}_{\varphi,1234,\text{int}}L$
(c) Graph of resultant $|R(\kappa)|$ for ${\cal Q}_{\varphi,34,\text{int}}L$
Fig 5: Graphs of the resultant $|R(\kappa)|$ for the idealized models of complex cells ${\cal Q}_{\varphi,12,\text{int}}L$ according to (11), ${\cal Q}_{\varphi,1234,\text{int}}L$ according to (12) and ${\cal Q}_{\varphi,34,\text{int}}L$ according to (13). (Horizontal axes in: scale parameter ratio $\kappa$. Vertical axes: resultant $|R|$.)
@@
Figs. 6a–c show corresponding resultant histograms for each of these integrated affine quasi quadrature measures, when assuming a uniform distribution over a logarithmic parameterization of the scale parameter ratio $\kappa$ over the interval $1/\kappa_{\max},\kappa_{\max}]$ for the again arbitrary value of $\kappa_{\max}=8$. Such a logarithmic distribution constitutes a natural default prior for a strictly positive variable according to Jaynes [101]. Fig. 6d shows a combined histogram of the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$, when assuming an equal number of idealized complex cells for these two types.
(a) Histogram over $|R|$ for ${\cal Q}_{\varphi,12,\text{int}}L$
(b) Histogram over $|R|$ for ${\cal Q}_{\varphi,1234,\text{int}}L$
(c) Histogram over $|R|$ for ${\cal Q}_{\varphi,34,\text{int}}L$
(d) Combined histogram over $|R|$ for ${\cal Q}_{\varphi,12,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$
Fig 6: Orientation selectivity histograms for the resultant $|R|$ for (a–c) the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$, ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ according to (11), (12) and (13), when assuming a uniform distribution over a logarithmic parameterization of the scale parameter ratio $\kappa$, as well as (d) combined histogram of ${\cal Q}_{\varphi,12,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$, when assuming equal numbers of complex cells for these two types. (Horizontal axes: 10 quantized bins over the resultant $|R|\in[0,1]$. Vertical axes: bin counts.)
@@
As can be seen from these histograms, the use of the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,12,\text{int}}L$, ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ leads to resultant histograms with different shapes than for pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{pt}}L$ previously studied in Lindeberg [14]. Specifically, except for the peak at the bin corresponding to $|R|\in[0.6,0.7]$, the resultant histogram of ${\cal Q}_{\varphi,1234,\text{int}}L$ is more uniform and without any bias the bias towards either smaller values of the resultant $|R|$ towards the bins $|R|\in[0.1,0.2]$ and $|R|\in[0.7,0.8]$ as for the resultant histogram of the pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{pt}}L$.
@@
Additionally, when including third- and fourth-order terms, the integrated affine quasi quadrature measures ${\cal Q}_{\varphi,1234,\text{int}}L$ and ${\cal Q}_{\varphi,34,\text{int}}L$ lead to accumulations to the bins $|R|\in[0.8,0.9]$ and $|R|\in[0.9,1.0]$, while those bins are not really reached by the pointwise quasi quadrature measure ${\cal Q}_{\varphi,12,\text{pt}}L$. In these respects, this analysis suggests that the mechanisms of spatial integration and inclusion of higher-order terms than mere first- and second-order terms may be important to more quantitatively model the orientation selectivity properties of complex cells. Furthermore, one may speculate if a suitable reformulation of the non-linearities in the composed integrated affine quasi quadrature measure ${\cal Q}_{\varphi,1234,\text{int}}L$ could move the peak for the bin $|R|\in[0.6,0.7]$ to the bin positions $|R|\in[0.3,0.4]$ and $|R|\in[0.5,0.6]$ that lead to peaks in the resultant histogram for the biological complex cells.